# InstructGIE: Towards Generalizable Image Editing

### v1

@@ Abstract
Recent advances in image editing have been driven by the development of denoising diffusion models, marking a significant leap forward in this field. Despite these advances, the generalization capabilities of recent image editing approaches remain constrained. In response to this challenge, our study introduces a novel image editing framework with enhanced generalization robustness by boosting in-context learning capability and unifying language instruction. This framework incorporates a module specifically optimized for image editing tasks, leveraging the VMamba Block and an editing-shift matching strategy to augment in-context learning. Furthermore, we unveil a selective area-matching technique specifically engineered to address and rectify corrupted details in generated images, such as human facial features, to further improve the quality. Another key innovation of our approach is the integration of a language unification technique, which aligns language embeddings with editing semantics to elevate the quality of image editing. Moreover, we compile the first dataset for image editing with visual prompts and editing instructions that could be used to enhance in-context capability. Trained on this dataset, our methodology not only achieves superior synthesis quality for trained tasks, but also demonstrates robust generalization capability across unseen vision tasks through tailored prompts.
@@ 1 Introduction
On the other hand, large language models (LLMs) have proven extraordinary abilities to learn from contexts, referred to as in-context learning, which allows LLMs to perform unseen tasks by providing the combination of input-output examples and a query input. Inspired by the potential to enhance the generalization of the model with LLMs, [38, 41] explore in-context learning for computer vision tasks, allowing them to perform unseen tasks with novel vision-language prompt designs. However, these methods are not tailored for image editing applications, leading to unsatisfying synthetic qualities with inaccurate or incorrect output and lack of detail. To improve the generalization of image editing with improved synthetic image quality, it is crucial to effectively understand the text & image prompts and specifically control image editing details, which is challenging in the current literature.
In this work, we propose InstructGIE, an image editing framework with enhanced generalizability. We improve image editing performance from both visual and text aspects. (i) For the visual information, we incorporate a VMamba-based module to specifically enhance the image editing outputs. As VMamba[26] has proven its better performance in capturing in-context information from inputs with larger receptive fields[26], we leverage VMamba and propose an editing-shift matching strategy to augment in-context learning. Furthermore, since current image editing works do not perform well to generate correct features with accurate details, we unveil a selective area-matching technique specifically engineered to address and rectify corrupted details in generated images, such as human facial features, to further improve the quality. (ii) Another key innovation of our approach is the integration of a language unification technique, which aligns language embeddings with editing semantics to elevate the quality of image editing. Our framework not only achieves superior in-context generation for trained tasks but also demonstrates robust generalization across unseen vision tasks. Moreover, we compile a publicly available image editing dataset with plenty of visual prompts and editing instructions for better generalization evaluation of image editing. Our contributions are summarized as follows:
@@ 2.3 Dataset for Diffusion-based Image Editing
Currently, various types of datasets are used for training in diffusion-based image editing. There are datasets that concentrate on specific domains like CelebA [27] and FFHQ [19] for human face image manipulation, AFHQ [7] for animal face image editing, LSUN [42] for object modification, and WikiArt [29] for style transfer. In-the-wild video datasets could also be leveraged to train image editing tasks. The Scannet dataset [9] encompasses a vast array of more than 1,500 indoor scenes from various settings such as apartments, offices, and hotels, providing extensive annotations. The LRW dataset [8], tailored for lip reading tasks, includes more than 1000 video utterances of 500 distinct words. The UBC-Fashion dataset [43] features 600 videos spanning various clothing categories, with 500 videos allocated for training and 100 for testing, guaranteeing no repetition of individuals in the training set. The DAVIS dataset[43] (Densely Annotated VIdeo Segmentation), a widely recognized benchmark for video object segmentation, contains 150 videos in total. There are also image editing works proposing to generate datasets with editing instructions. InstructPix2pix [3] collects over 450,000 training image pairs. For each pair, given an image with its caption, it first uses a finetuned GPT-3 [4] to generate an editing instruction and an edited image caption. Then it employs Stable Diffusion and the Prompt-to-Prompt algorithm [14] to generate edited image following the caption. However, currently there are no datasets with multiple image pairs under one editing instruction, which is crucial to enhance the generalization ability of image editing.
@@ 3 Preliminary
Expanding this framework to conditional generative modeling, the process involves conditioning the neural network $f_{\theta}\left(\mathbf{x}_{t},t,\mathbf{y}\right)$ on an additional input $\mathbf{y}$, enabling the generation of images from a distribution conditioned on $\mathbf{y}$. This conditional input could be anything from a low-resolution image, a category label, to a descriptive text sequence. Leveraging the advancements in LLMs [33] and hybrid vision-language models [31], text-to-image diffusion models are developed. These models allow for the creation of detailed, high-resolution images from mere text descriptions, starting with a low-resolution image generated through the diffusion process, which is subsequently refined into a high-resolution image using additional models.
@@ 4 The Proposed Method
Figure 2: Overall architecture of our proposed framework. The upper pipepline is used for training only, and the lower pipeline is for both training and inference process. During training, the output image from the lower pipeline is fed into the upper pipeline to obtain the losses.
@@ 4.1 Enhanced In-Context Learning
To improve the generalization ability of image editing, it is crucial for the framework to explicitly capture low-level visual editing contexts. Current diffusion-based image editing methods [38, 30] that involve visual prompting either adopt ConvNet [23] or ViT [11] as the vision encoder for visual prompts. However, these methods fail to generalize well as they are not able to capture enough visual editing contexts (see Fig. 3). To address this, we formulate the visual prompted condition as a single image $\textbf{Img}^{\text{VPcon}}$ = $\textbf{Img}_{0}^{\text{in}},$ $\textbf{Img}_{0}^{\text{out}},$ $\textbf{Img}_{1}^{\text{in}},$ Grey}, as shown in Fig 2, with a global effective receptive field (ERF) to better capture visual contexts.
Moreover, we further propose to reform the conditioned latent diffusion model. Inspired by the recent visual state space model VMamba [26] which exhibits a better global ERF, we propose to adopt a vision encoder based on Zero-VMamba to fit our structure. Specifically, the vision encoder $\mathcal{G}_{ven}$ comprehends the visual prompted condition in latent space $x_{vpc}$ as follows,
@@
To further improve the performance, after each trainable copied modules $\mathcal{F}_{c}$ with parameters $\Theta_{c}$, we propose to link and inject the processed visual prompted condition information to the frozen Stable Diffusion model $\mathcal{F}$ with parameters $\Theta$ through zero-VMamba layer $\mathcal{G}$. We use two instances of VMamba with parameters $\theta_{g1}$ and $\theta_{g2}$ respectively. The complete model then computes the following,
@@ 4.2 Language Instruction Unification
We then augment the training data with these newly unified language editing instructions.
@@ 5 Visual Prompt Dataset For Image Editing
Therefore, we introduce and open source a new dataset that is designed specifically for image editing tasks utilizing visual prompts. Our dataset generation pipeline is shown in Fig. 4 with 2 phases: Data Generation and Data Processing. In Data Generation phase, we first fine-tune GPT-3 [4] for 2 epochs with 700 groups of human-written edits, each consisting of 5 different pairs of input and edited captions with one editing instruction. Then as shown in step 1, we generate around 3500 groups of editing prompts using the fine-tuned GPT-3 model. In each group, there is one instruction $E_{0}$ and five pairs of caption and edited caption $(I_{0},O_{0}),(I_{1},O_{1}),...,(I_{4},O_{4})$. In step 2, similar to InstructPix2Pix, we then also adopt Prompt-to-Prompt for image generation. For each input and edited caption pair, we generated 50 times each with random noise and followed InstructPix2Pix to keep the best one image pair using CLIP-based metric. In addition, we also make sure for each editing instruction, there are at least two pairs of images. In step 3, we generate more image pair sets using in the same way. We generate around 10,000 image pairs with around 3,000 editing instructions. In Data Preparing phase, we randomly choose 2 pairs of images under the same editing instruction and concatenate them together to get one image for training.
@@ 6.1 Experimental Settings
Datasets. To fairly compare our methods with baseline methods, we conduct both qualitative and quantitative experiments on our proposed synthetic dataset that involves over 10,000 image pairs and 3,000 editing instructions. All single input images have a resolution of $512\times 512$ and are resized to $256\times 256$ for training and testing purposes.
@@ 7 Conclusion
In this work, we propose InstructGIE, an image editing framework with enhanced generalization ability. We improve the image editing performance from both visual and text aspects. We incorporate a module based on VMamba to specifically enhance the image editing outputs for the visual information. We further propose an editing-shift matching strategy to augment in-context learning. Furthermore, we unveil a selective area-matching technique specifically engineered to address and rectify corrupted details in generated images, such as human facial features, to further improve the quality. Another key innovation of our approach is the integration of a language unification technique, which aligns language embeddings with editing semantics to elevate the quality of image editing. Moreover, we compile a publicly available image editing dataset with plenty of visual prompts and editing instructions for better generalization evaluation of image editing. Extensive experiments has demonstrated that our framework not only achieves superior in-context generation for trained tasks, but also demonstrates robust generalization across unseen vision tasks, both quantative and qualitively.

### v2

@@ Abstract
Recent advances in image editing have been driven by the development of denoising diffusion models, marking a significant leap forward in this field. Despite these advances, the generalization capabilities of recent image editing approaches remain constrained. In response to this challenge, our study introduces a novel image editing framework with enhanced generalization robustness by boosting in-context learning capability and unifying language instruction. This framework incorporates a module specifically optimized for image editing tasks, leveraging the VMamba block and an editing-shift matching strategy to augment in-context learning. Furthermore, we unveil a selective area-matching technique specifically engineered to address and rectify corrupted details in generated images, such as human facial features, to further improve the quality. Another key innovation of our approach is the integration of a language unification technique, which aligns language embeddings with editing semantics to elevate the quality of image editing. Moreover, we compile the first dataset for image editing with visual prompts and editing instructions that could be used to enhance in-context capability. Trained on this dataset, our methodology not only achieves superior synthesis quality for trained tasks, but also demonstrates robust generalization capability across unseen vision tasks through tailored prompts. Our project page is available at https://cr8br0ze.github.io/InstructGIE.
@@ 1 Introduction
On the other hand, large language models (LLMs) have proven extraordinary abilities to learn from contexts, referred to as in-context learning, which allows LLMs to perform unseen tasks by providing a combination of input-output examples and a query input. Inspired by the potential to enhance the generalization of the model with LLMs, [38, 41] explore in-context learning for computer vision tasks, allowing them to perform unseen tasks with novel vision-language prompt designs. However, these methods are not tailored for image editing applications, leading to unsatisfying synthetic qualities with inaccurate or incorrect output and lack of detail. To improve the generalization of image editing with improved synthetic image quality, it is crucial to effectively understand the text & image prompts and specifically control image editing details, which is challenging in the current literature.
In this work, we propose InstructGIE, an image editing framework with enhanced generalizability. We improve image editing performance from both visual and text aspects. (i) For the visual information, we incorporate a VMamba-based module to specifically enhance the image editing outputs. As VMamba[26] has proven its better performance in capturing in-context information from inputs with larger receptive fields[26], we leverage VMamba and propose an editing-shift matching strategy to augment in-context learning. Furthermore, since current image editing works do not perform well in generating correct features with accurate details, we unveil a selective area-matching technique specifically engineered to address and rectify corrupted details in generated images, such as human facial features, to further improve the quality. (ii) Another key innovation of our approach is the integration of a language unification technique, which aligns language embeddings with editing semantics to elevate the quality of image editing. Our framework not only achieves superior in-context generation for trained tasks but also demonstrates robust generalization across unseen vision tasks. Moreover, we compile a publicly available image editing dataset with plenty of visual prompts and editing instructions for better generalization evaluation of image editing. Our contributions are summarized as follows:
@@ 2.3 Dataset for Diffusion-based Image Editing
Currently, various types of datasets are used for training in diffusion-based image editing. There are datasets that concentrate on specific domains like CelebA [27] and FFHQ [19] for human face image manipulation, AFHQ [7] for animal face image editing, LSUN [42] for object modification, and WikiArt [29] for style transfer. In-the-wild video datasets could also be leveraged to train image editing tasks. The Scannet dataset [9] encompasses a vast array of more than 1,500 indoor scenes from various settings, such as apartments, offices, and hotels, providing extensive annotations. The LRW dataset [8], tailored for lip reading tasks, includes more than 1000 video utterances of 500 distinct words. The UBC-Fashion dataset [43] features 600 videos spanning various clothing categories, with 500 videos allocated for training and 100 for testing, guaranteeing no repetition of individuals in the training set. The DAVIS dataset[43] (Densely Annotated VIdeo Segmentation), a widely recognized benchmark for video object segmentation, contains 150 videos in total. There are also image editing works proposing to generate datasets with editing instructions. InstructPix2pix [3] collects over 450,000 training image pairs. For each pair, given an image with its caption, it first uses a finetuned GPT-3 [4] to generate an editing instruction and an edited image caption. Then it employs Stable Diffusion and the Prompt-to-Prompt algorithm [14] to generate edited image following the caption. However, currently there are no datasets with multiple image pairs under one editing instruction, which is crucial to enhance the generalization ability of image editing.
@@ 3 Preliminary
Expanding this framework to conditional generative modeling, the process involves conditioning the neural network $f_{\theta}\left(\mathbf{x}_{t},t,\mathbf{y}\right)$ on an additional input $\mathbf{y}$, enabling the generation of images from a distribution conditioned on $\mathbf{y}$. This conditional input could be anything from a low-resolution image, a category label, or a descriptive text sequence. Leveraging the advancements in LLMs [33] and hybrid vision-language models [31], text-to-image diffusion models are developed. These models allow for the creation of detailed, high-resolution images from mere text descriptions, starting with a low-resolution image generated through the diffusion process, which is subsequently refined into a high-resolution image using additional models.
@@ 4 The Proposed Method
Figure 2: Overall architecture of InstructGIE. The lower pipeline is for both training and inference processes where the model obtains unified editing instructions outputted by Instruction Unification Module $\mathcal{U}$ and combines with visual prompted input $\textbf{Img}^{\text{VPcon}}$ to pass through Zero-VMamba integrated Stable Diffusion model with ControlNet for output image. The upper pipeline is for training only which compares output image and training ground truth $\textbf{Img}^{\text{train}}$ and computes editing shift loss $\mathcal{L}_{es}$ with Editing Shift Module and selective area matching loss $\mathcal{L}_{sam}$ with Selective Area Matching Module.
@@ 4.1 Enhanced In-Context Learning
To improve the generalization ability of image editing, it is crucial for the framework to explicitly capture low-level visual editing contexts. Current diffusion-based image editing methods [38, 30] that involve visual prompting either adopt ConvNet [23] or ViT [11] as the vision encoder for visual prompts. However, these methods fail to generalize well as they are not able to capture enough visual editing contexts (see Fig. 3). To address this, we formulate the visual prompted condition as a single image $\textbf{Img}^{\text{VPcon}}$ = $\textbf{Img}_{0}^{\text{in}},$ $\textbf{Img}_{0}^{\text{out}},$ $\textbf{Img}_{1}^{\text{in}},$ Grey}, as shown in Fig 2, with a global effective receptive field (ERF).
Moreover, we further propose to reform the conditioned latent diffusion model. Inspired by the recent visual state space model VMamba [26] which exhibits a better global ERF and also emphasizes shifting boundaries of input’s four quadrants, we propose to adopt a vision encoder based on Zero-VMamba to fit our structure. Specifically, the vision encoder $\mathcal{G}_{ven}$ comprehends the visual prompted condition in latent space $x_{vpc}$ as follows,
@@
To further improve the performance, after each ControlNet trainable copied modules $\mathcal{F}_{c}$ with parameters $\Theta_{c}$, we propose to link and inject the processed visual prompted condition information to the frozen Stable Diffusion model $\mathcal{F}$ with parameters $\Theta$ through zero-VMamba layer $\mathcal{G}$. We use two instances of VMamba with parameters $\theta_{g1}$ and $\theta_{g2}$ respectively. The complete model then computes the following,
@@ 4.2 Language Instruction Unification
We then augment the training data with unified language editing instructions.
@@ 5 Visual Prompt Dataset For Image Editing
Therefore, we introduce and open source a new dataset that is designed specifically for image editing tasks utilizing visual prompts. Our dataset generation pipeline is shown in Fig. 4 with 2 phases: Data Generation and Data Processing. In Data Generation phase, we first fine-tune GPT-3 [4] for 2 epochs with 700 groups of human-written edits, each consisting of 5 different pairs of input and edited captions with one editing instruction. Then as shown in step 1, we generate around 3500 groups of editing prompts using the fine-tuned GPT-3 model. In each group, there is one instruction $E_{0}$ and five pairs of caption and edited caption $(I_{0},O_{0}),(I_{1},O_{1}),...,(I_{4},O_{4})$. In step 2, similar to InstructPix2Pix, we then also adopt Prompt-to-Prompt for image generation. For each input and edited caption pair, we generated 50 times each with random noise and followed InstructPix2Pix to keep the best one image pair using CLIP-based metric. In addition, we also make sure for each editing instruction, there are at least two pairs of images. In step 3, we generate more image pair sets using in the same way. With filtering, we obtained around 12,000 images with around 2,000 editing instructions. In data preparing, we randomly choose 2 pairs of images under the same editing instruction and concatenate them for training.
@@ 6.1 Experimental Settings
Datasets. To fairly compare our methods with baseline methods, we conduct both qualitative and quantitative experiments on our proposed synthetic dataset that involves over 12,000 filtered images and 2,000 filtered editing instructions. All single input images have a resolution of $512\times 512$ and are resized to $256\times 256$ for training and testing purposes.
@@ 7 Conclusion
In this work, we propose InstructGIE, an image editing framework with enhanced generalization ability, improving performance in both visual and text aspects. We incorporate a VMamba-based module to enhance visual outputs and introduce an editing-shift matching strategy to augment in-context learning. Our selective area-matching technique addresses and rectifies corrupted details, while a language unification technique aligns language embeddings with editing semantics. Additionally, we compile a publicly available dataset for better generalization evaluation. Extensive experiments demonstrate our framework’s superior in-context generation performance and robust generalization capability across unseen vision tasks, both quantative and qualitively.
Acknowledgement
This work is partially supported by the Army Research Office/Army Research Laboratory via grant W911-NF-20-1-0167 to Northeastern University, National Science Foundation CCF-1937500, and the Fundamental Research Funds for the Central Universities, Peking University.